# Fluid mechanics, Applied mathematics, Mathematical modelling

### v2

@@ 6 Conclusion
Appendix
Appendix A Full model specification
In this appendix we provide a complete statement of the model. The model is given by the following equations for the six fields $u_{0},u_{1},v_{1},w_{0},w_{1},q_{0}$, each depending on streamwise coordinate $x$, spanwise coordinate $z$, and time $t$,
$\frac{\partial u_{0}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})u_{0}+\frac{1}{2}(u_{1}\partial_{x}+\beta v_{1}+w_{1}\partial_{z})u_{1}=-\partial_{x}p_{0}+\left(\frac{1}{Re}\nabla^{2}_{H}-\alpha_{0}\right)u_{0},$
$\frac{\partial u_{1}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})u_{1}+(u_{1}\partial_{x}+w_{1}\partial_{z})u_{0}=-\partial_{x}p_{1}+\left(\frac{1}{Re}(\nabla^{2}_{H}-\beta^{2})-\alpha\right)u_{1}+f-\uwave{\beta A(q_{0})},$
$\frac{\partial v_{1}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})v_{1}=-\beta p_{1}+\frac{1}{Re}(\nabla^{2}_{H}-\beta^{2})v_{1}+\uwave{\partial_{x}A(q_{0})},$
$\frac{\partial w_{0}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})w_{0}+\frac{1}{2}(u_{1}\partial_{x}+\beta v_{1}+w_{1}\partial_{z})w_{1}=-\partial_{z}p_{0}+\left(\frac{1}{Re}\nabla^{2}_{H}-\alpha_{0}\right)w_{0},$
$\frac{\partial w_{1}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})w_{1}+(u_{1}\partial_{x}+w_{1}\partial_{z})w_{0}=-\partial_{z}p_{1}+\left(\frac{1}{Re}(\nabla^{2}_{H}-\beta^{2})-\alpha\right)w_{1},$
$=\frac{1}{2}\uwave{A(q_{0})}[\beta u_{1}+\partial_{x}v_{1}]-2\alpha q_{0}-\uwave{\varepsilon_{0}(q_{0};Re)}+\frac{1}{Re}\nabla^{2}_{H}q_{0}+\uwave{\nabla_{H}\cdot\left(\nu_{T}(q_{0};Re)\nabla_{H}q_{0}\right)},$
where $f=\alpha+\beta^{2}/Re$, $\beta=\pi/2$, and $\nabla_{H}=(\partial_{x},\partial_{z})$. The terms coming from the turbulence closure and highlighted by wavy underline. The only forcing of the large-scale flow by Reynolds stresses appears only via two terms. All other terms in the momenta equations arises just from modelling by Ekman friction and a Galerkin projection of the Navier-Stokes equations onto the retained modes.
The equations are closed with the following,
Equations (35) constitute the turbulence closures, expressing the dependence of Reynold stress, pseudo-dissipation, and turbulent transport on turbulence amplitude $q_{0}$. Equation (36) follows from the quasi-static approximation for the $q_{1}$ field.
Model parameters have the values
$a=0.3,\,c=9.65,\,d=0.05,\,\eta=5\times 10^{-3},\,\alpha=0.01,\,\alpha_{0}=0.03,\,\kappa=0.09.$ (37)
@@ Appendix B Linear-stability equations
For a detailed derivation of the linear-stability equations, see the electronic supplementary material [55].

### v3

@@ 6 Conclusion
Appendix A Model equation in a 1D tilted domain
In this appendix we present the model equations restricted to solutions at a prescribed angle $\theta$ with respect to the streamwise direction. For this purpose, we work in horizontal coordinates $({\bar{x}},{\bar{z}})$ rotated by angle $\theta$ with respect to the streamwise-spanwise coordinates $(x,z)$. The bases vectors, coordinates, and horizontal velocity components obey
${\bf e}_{\bar{x}}$ $=\cos\theta\,{\bf e}_{x}-\sin\theta\,{\bf e}_{z},$ ${\bf e}_{\bar{z}}=\sin\theta\,{\bf e}_{x}+\cos\theta\,{\bf e}_{z}$ (39)
${\bar{x}}$ $=\cos\theta\,x+\sin\theta\,z,$ ${\bar{z}}=-\sin\theta x+\cos\theta\,z,$ (40)
${\bar{u}}$ $=\cos\theta\,u-\sin\theta\,w,$ ${\bar{w}}=\sin\theta\,u+\cos\theta\,w,$ (41)
where rotated quantities appear on the left. We take ${\bf e}_{\bar{x}}$ to point in the along-band direction and ${\bf e}_{\bar{z}}$ in the across-band direction. The wall-normal coordinate $y$ and velocity $v$, and scalar TKE $q$ are unaffected by horizontal rotation. The model simplification comes by considering solutions that are independent of the along-band direction $\bar{x}$}. Hence, model fields become functions of $({\bar{z}},t)$ only and all along-band derivatives, $\partial_{{\bar{x}}}$, vanish from the governing equations.
While applying the change of coordinates is straightforward, the system is not isotropic and some terms in the governing equations explicitly depend on streamwise direction ${\bf e}_{x}$. These terms pick up $\cos\theta$ or $\sin\theta$ factors when expressed in rotated coordinates. Specifically, the driving body force takes the form
${\bf f}=(\alpha+\beta^{2}/Re)\sin\left(\beta y\right){\bf e}_{x}=(\alpha+\beta^{2}/Re)\sin\left(\beta y\right)(\cos\theta\,{\bf e}_{\bar{x}}+\sin\theta\,{\bf e}_{\bar{z}})$
The divergence of the Reynolds stress takes the form
$\nabla\cdot\mathcal{R}$ $=-\beta A(q_{0})\sin(\beta y){\bf e}_{x}+\partial_{x}A(q_{0})\cos(\beta y){\bf e}_{y}$
$=-\beta A(q_{0})\sin(\beta y)(\cos\theta\,{\bf e}_{\bar{x}}+\sin\theta\,{\bf e}_{\bar{z}})+\sin\theta\,\partial_{\bar{z}}A(q_{0})\cos(\beta y){\bf e}_{y},$
where in the last term we have used that $\partial_{x}=\cos\theta\,\partial_{{\bar{x}}}+\sin\theta\,\partial_{{\bar{z}}}=\sin\theta\,\partial_{{\bar{z}}}$. Finally, the production terms take the form
$\mathcal{P}=(\partial_{y}u+\partial_{x}v)A(q_{0})\cos(\beta y)=(\cos\theta\,\partial_{y}{\bar{u}}+\sin\theta\,\partial_{y}{\bar{w}}+\sin\theta\,\partial_{\bar{z}}v)A(q_{0})\cos(\beta y)$
Thus, the model equations in rotated coordinates become
$\frac{\partial{\bar{u}}_{0}}{\partial t}$ $+\frac{1}{2}(\beta v_{1}+{\bar{w}}_{1}\partial_{{\bar{z}}}){\bar{u}}_{1}=\left(\frac{1}{Re}\partial_{{\bar{z}}}^{2}-\alpha_{0}\right){\bar{u}}_{0},$
$\frac{\partial{\bar{u}}_{1}}{\partial t}$ $+{\bar{w}}_{1}\partial_{{\bar{z}}}{\bar{u}}_{0}=\left(\frac{1}{Re}(\partial_{{\bar{z}}}^{2}-\beta^{2})-\alpha\right){\bar{u}}_{1}+f\cos\theta-\beta A(q_{0})\cos\theta,$
$\frac{\partial v_{1}}{\partial t}$ $=-\beta p_{1}+\frac{1}{Re}(\partial_{{\bar{z}}}^{2}-\beta^{2})v_{1}-\partial_{{\bar{z}}}A(q_{0})\sin\theta,$
$\frac{\partial{\bar{w}}_{1}}{\partial t}$ $=-\partial_{{\bar{z}}}p_{1}+\left(\frac{1}{Re}(\partial_{{\bar{z}}}^{2}-\beta^{2})-\alpha\right){\bar{w}}_{1}+\beta A(q_{0})\sin\theta-f\sin\theta,$
$\frac{\partial q_{0}}{\partial t}$ $+\frac{1}{2}(\beta v_{1}+{\bar{w}}_{1}\partial_{{\bar{z}}})q_{1}({\bar{u}}_{1},{\bar{w}}_{1},q_{0})=\frac{1}{2}A(q_{0})[\beta({\bar{u}}_{1}\cos\theta-{\bar{w}}_{1}\sin\theta)-\partial_{{\bar{z}}}v_{1}\sin\theta]$
$-2\alpha q_{0}-\varepsilon_{0}(q_{0};Re)+\frac{1}{Re}\partial_{{\bar{z}}}^{2}q_{0}+\partial_{{\bar{z}}}\left(\nu_{T}(q_{0};Re)\partial_{{\bar{z}}}q_{0}\right),$
where $f=\alpha+\beta^{2}/Re$ and $\beta=\pi/2$. The fields ${\bar{u}}_{0}({\bar{z}},t)$ and ${\bar{u}}_{1}({\bar{z}},t)$ capture flow along the bands, while ${\bar{w}}_{1}({\bar{z}},t)$ captures flow across the bands. The field ${\bar{w}}_{0}$, describing the vertically constant across-band flow, vanishes by incompressibility: $\partial_{{\bar{x}}}{\bar{u}}_{0}+\partial_{{\bar{z}}}{\bar{w}}_{0}=\partial_{{\bar{z}}}{\bar{w}}_{0}=0$, from which ${\bar{w}}_{0}=0$, assuming no bulk flow. The system of equations can be further reduced to only four independent fields, since incompressibility of the first vertical modes implies $\beta v_{1}=\partial_{{\bar{z}}}{\bar{w}}_{1}$. Two possible approaches are: replace the evolution equations for the $v_{1}$ and ${\bar{w}}_{1}$ fields with a single evolution equation for the vorticity component $\zeta_{1}=\beta{\bar{w}}_{1}-\partial_{\bar{z}}v_{1}$, or use incompressibility to obtain an expression for the pressure $p_{1}$. The two approaches are ultimately equivalent and result in the simplest formulation of the 1D model equations. We have solved these equations numerically to produce figure 8.
@@ Appendix B Linear-stability equations
For a detailed derivation of the linear-stability equations, see the electronic supplementary material [SM].